# Sparse Portfolio Selection via Topological Data Analysis based Clustering

### v1

@@ Abstract
This paper uses topological data analysis (TDA) tools and introduces a data-driven clustering-based stock selection strategy tailored for sparse portfolio construction. Our asset selection strategy exploits the topological features of stock price movements to select a subset of topologically similar (different) assets for a sparse index tracking (Markowitz) portfolio. We introduce new distance measures, which serve as an input to the clustering algorithm, on the space of persistence diagrams and landscapes that consider the time component of a time series. We conduct an empirical analysis on the S&P index from 2009 to 2020, including a study on the COVID-19 data to validate the robustness of our methodology. Our strategy to integrate TDA with the clustering algorithm significantly enhanced the performance of sparse portfolios across various performance measures in diverse market scenarios.
keywords:
Topological Data Analysis, Clustering, Index Tracking, Mean-Variance Portfolio, Global Minimum Variance Portfolio, Sparse Portfolios
journal: Decision Support Systems
@@ 1 Introduction
Asset selection and portfolio optimization are critical tasks for financial institutions such as banks, fund management firms, and individual investors as they seek to manage risk and maximize returns in an increasingly competitive business environment and complex market conditions. Over the years, academics and practitioners have actively developed quantitative techniques and models to assess risk and construct efficient portfolios that perform well out-of-sample. In this pursuit, sparse portfolio selection has become a well-known and extensively researched problem in the fields of financial economics and optimization, where sparseness is a desirable property not only from the point of view of transaction costs but also because it allows investors to concentrate their capital on assets that are most likely to generate returns while minimizing exposure to unforeseen risks or market fluctuations. Moreover, it also offers a potential solution to the limitations of traditional mean-variance or index-tracking portfolios, as it focuses on a smaller subset of assets to improve portfolio performance. [^1] Despite its popularity, mean-variance analysis has received criticism for its susceptibility to parameter estimates and assumptions. It is observed that small changes in assumed asset returns, volatilities, or correlations lead to substantial effects on the optimization results, ultimately leading to sub-optimal performance when tested against real-world data. In essence, the classic Markowitz portfolio optimization represents an ill-posed inverse problem. Several techniques have been proposed to address this challenge, including regularization techniques imposing sparseness and adopting shrinkage estimators of expected returns and covariance matrix. [^2] In index tracking, the goal is to construct a portfolio of securities that replicates the performance of a market index. The simplest method to achieve this goal is through full replication, which entails holding all the assets in the target market index in the same proportions. While full replication provides perfect tracking performance in a frictionless market, it is expensive, requiring frequent re-balancing of many stocks. It is sometimes not possible due to restrictions on trading illiquid securities or managerial constraints. An alternative approach is partial replication, which aims to mimic the index performance using a small subset of securities.
@@
The problem of sparse portfolio selection can be reformulated as a two-step problem, with the first step of selecting a subset of securities based on their risk and return characteristics and the second step of obtaining the optimal weights. In this paper, we adopt clustering to select a smaller set of securities that provide diversification benefits to the portfolio. Clustering methods have emerged as one of the most innovative strategies in portfolio selection in recent years. This technique for classifying data objects into homogeneous groups based on their similarities proved to be effective in constructing sparse portfolios (Panton et al. (1976); Hui (2005)).
@@
This two-step approach has several advantages over the alternative methods in the literature, such as regularization techniques (imposing a constraint on the $\ell_{q}$-norm of portfolio weights) or including a cardinality constraint on the number of securities to be included in the portfolio to promote sparsity and stable allocations. Adding a cardinality constraint significantly increases the combinatorial complexity of the optimization problem, resulting in a mixed-integer quadratic programming problem that is generally NP-hard and computationally intractable. Heuristic methods, such as genetic algorithms, are typically required to find approximate solutions, which may not guarantee global optimality (Ruiz-Torrubiano and Suárez (2009); Woodside-Oriakhi et al. (2011); Fastrich et al. (2014)). Similarly, the regularization technique involves determining the regularization coefficients through cross-validation, which becomes particularly challenging when each security has its regularization parameter. This approach has been shown to outperform the approach with a common regularization parameter. [^3] For example, Brodie et al. (2009) suggests integrating an $\ell_{1}$-norm (LASSO) penalty on the vector of portfolio weights in optimization models to regularize the optimization problem and encourage sparse portfolios. Moreover, DeMiguel et al. (2009a) demonstrate that norm-constrained global minimum variance portfolio yields improved out-of-sample performance and Giamouridis and Paterlini (2010) and Giuzio et al. (2018) utilize $\ell_{1}$ penalty in the context of replicating hedge funds. Recently, Ho et al. (2015) considered the mean-variance optimization with an advanced regularization technique called weighted Elastic-Net. Further, these methods tend to result in risk concentration as the resulting portfolio may not be diversified. This can lead to over-exposure to a single industry sector, thus making the portfolio riskier due to vulnerability to a downturn in that sector. Furthermore, due to local optima and discontinuities in the search space, both the cardinality constraint problem and the $\ell_{q}$-norm constraint problem are difficult to solve from an optimization standpoint. Finally, the regularization technique LASSO suffers from several known shortcomings, such as selecting at random from equally correlated assets (Bondell and Reich (2008)), being biased for large coefficient values (Fan and Li (2001); Giuzio et al. (2018)), and being ineffective in the no-short-sale area (DeMiguel et al. (2009a)). In contrast, clustering can result in a more diversified and sparse portfolio without increasing the computational burden. Moreover, clustering allows for the combination of judgment and quantitative optimization as it is an intuitive approach.
@@
The key to any clustering technique is the distance function that enables it to identify similarities and dissimilarities among the inputs in the data set (time series of stock returns in our case) and allocate them to different clusters. Traditionally, correlation-based clustering has been employed to locate clusters in the portfolio of financial assets, for instance, Mantegna (1999). However, it is well known that Pearson’s correlation is not a suitable measure of dependence to be used outside the Gaussian (elliptical) distributions, and, in particular, it does not give an accurate indication and understanding of the actual dependence between risk exposures Embrechts et al. (2003). Furthermore, even if two assets are perfectly correlated, an investment in one can lead to positive and negative returns in the other. In this paper, we employ a recently introduced approach known as Topological Data Analysis (TDA) to define the similarity between two time series and propose an efficient data-driven approach to sparse portfolio selection. The underlying idea of TDA is that persistent homology, through filtration, enables highly flexible and multi-resolution analysis of $n$-dimensional holes, which are the most salient features of a data set. TDA thus allows us to relate objects based on their qualitative structural properties, such as shape and connectedness. This paper argues that topological features can help identify exciting time series patterns and spatial data clustering. In particular, insights can be gained by examining the essential qualitative characteristics of data sets that are often overlooked by traditional distance-based clustering techniques. To our knowledge, this paper is the first to present a pioneering application of TDA to address the sparse portfolio selection using clustering.
@@
Our paper makes several methodological and empirical contributions that address the aforementioned issues. First, despite the considerable amount of research that has been conducted on applying TDA to study financial time-series, there still needs to be practical implementation and utilization of this method in the field of quantitative finance. [^4] Recently, Goel et al (2023) Goel et al. (2023) proposed TDA based sparse index tracking model where the authors adopt the regularization approach to achieve sparseness. We leverage tools from TDA, namely persistence diagrams and landscapes, and design strategies for sparse portfolio selection. By doing so, we establish a link between the growing research on TDA and the most frequently studied portfolio selection problem in the financial economics literature. Thus, we contribute to the emerging literature on applications of TDA to the problem of selecting sparse portfolios.
@@
We conduct an extensive empirical analysis, investigating the performance of the sparse index tracking, MV, and GMV portfolios obtained using the daily returns of the S&P 500 index and its constituents spanning approximately a decade from Dec 2009 to Dec 2019. We further assess the performance of our strategies during the recent COVID-19 Pandemic period from July 2019 to November 2020 to check the robustness. We also compare the performance of our sparse index tracking model on the benchmark data set for index tracking problems available publicly on the Beasley OR library. Our analysis shows that our TDA-based sparse portfolios outperform the traditional correlation-based sparse portfolios as well as other leading benchmark models in the literature.
@@ 2 Background of TDA
Definition 2.2.
The $p$-Wasserstein distance Cohen-Steiner et al. (2010) between two persistence diagrams ${\mathcal{D}}_{{X}}$ and ${\mathcal{D}}_{{Y}}$ corresponding to data sets $X$ and $Y$ is defined as
@@
Though persistence diagrams contain potentially valuable information about the underlying data set, it is a multi-set which, when equipped with Wasserstein distance, forms an incomplete metric space and hence is not appropriate to apply tools from statistics and machine learning for processing topological features (Bubenik et al. (2015)). A persistence landscape is an alternative tool to convey the information in the persistence diagram. It comprises a sequence of continuous and piece-wise linear functions defined on a re-scaled birth-death coordinate, with the peaks of a persistence landscape representing the significant topological features. Given the birth-death pairs $p_{i}(a_{i},b_{i}),~{}i=1,\ldots,m$, a persistence landscape, illustrated in Figure 3(b), is defined as a sequence of functions $\lambda:\mathbb{N}\times\mathbb{R}\rightarrow[0,\infty)$, denoted as $\lambda(k,t)=\lambda_{k}(t)$. Here, $\lambda_{k}(t)$ represents the $k$-th largest value among the set $\{\Lambda_{p_{i}}(t)\mid i=1,\ldots,m\}$ and we set $\lambda_{k}(x)=0$, ensuring that $\lambda_{k}(t)=0$ for $k>m$ if the $k$-th largest value does not exist. Here, for a fixed feature dimension $f$, we associate a piece-wise linear function $\Lambda_{p}(t):\mathbb{R}\rightarrow[0,\infty)$ with each birth-death pair $p=(a,b)\in{\mathcal{D}}$, where ${\mathcal{D}}$ is the persistence diagram, as follows:
@@
One of the notable advantages of the persistence landscape representation, as compared to persistence diagrams (refer to Bubenik et al. (2015); Bubenik (2018)), is that it forms a subset of the Banach space $L^{p}(\mathbb{N}\times\mathbb{R})$, which comprises sequences $\lambda=(\lambda_{k})_{k\in\mathbb{N}}$. This set possesses an inherent vector space structure and becomes a Banach space when equipped with a norm that measures the degree of “spread out” or “concentration” exhibited by the landscape functions. The norm of a persistence landscape characterizes a point cloud by a single numerical value, and a higher norm value indicates the presence of more significant topological features in the underlying point cloud and is given by
@@
Given that TDA operates on point clouds in multi-dimensional spaces, and our focus in this paper is on time series of stock returns, which lack natural point cloud representations, we must devise a method to obtain a point cloud from a time series. To achieve this, we adopt the standard procedure, called Taken’s embedding (Takens (1981)) that embeds a time series $x=\{x_{t},t=1,\ldots,T\}$ to a multi-dimensional representation as follows:
@@ 3 New distance measures
The notion of distance is inherently linked to the characteristics of the underlying data. For instance, the distance between two financial time series may differ from that between two DNA sequences. Nonetheless, distance measure aims to capture some sense of similarity, and often, one can employ the distance measure inherited from the reference space. However, if that is not the case, one must use a more suitable metric. For example, to compare persistence diagrams, Wasserstein distance (or bottleneck distance) is traditionally used due to their stability concerning perturbations of the input Gidea and Katz (2018); Gidea (2017); Goel et al. (2020). However, considering the time factor while dealing with financial time series is crucial, which the traditional Wasserstein distance fails to capture, as shown in the following toy example.
@@
While the example above is hypothetical, it advocates for a more appropriate distance measure for time-series data. It emphasizes that Wasserstein distance might be beneficial when dealing with observations from a fixed geometrical object where one is trying to identify the structure of that object. Still, it is unsuitable for financial applications where time is essential. Further, the literature has observed that computing the Bottleneck or the Wasserstein distances is computationally demanding as it requires finding point correspondence as observed in Agami (2023). Therefore, we propose new distance measures between persistence diagrams and persistence landscapes that are efficient in terms of time and also consider the time component inherent in time-series data.
@@
where ${\mathcal{D}}_{X-Y}$ represents the persistence diagram corresponding to the point cloud obtained from the series ${x}-{y}$ and ${\mathcal{D}}_{0}$ is the diagram corresponding to a constant series (note that PDs are shift-invariant), containing only the diagonal Mileyko et al. (2011), so that $WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{0})=\left(2^{-p}\sum_{a\in{\mathcal{D}}_{X}}(pers(a))^{p}\right)^{\frac{1}{p}}$, where $pers(a)$ denotes the persistence of $a=(b,d)\in{\mathcal{D}}_{x}$, given by $pers(a)=d-b$, with $b$ and $d$ denoting the birth and death time.

### v2

# Sparse Portfolio Selection via Topological Data Analysis-Based Clustering

@@ Abstract
This paper uses topological data analysis (TDA) tools and introduces a data-driven clustering-based stock selection strategy tailored for sparse portfolio construction. Our asset selection strategy exploits the topological features of stock price movements to select a subset of topologically similar (different) assets for a sparse index tracking (Markowitz) portfolio. We introduce new distance measures, which serve as an input to the clustering algorithm, on the space of persistence diagrams and landscapes that consider the time component of a time series. We conduct an empirical analysis on the S&P index from 2009 to 2022, including a study on the COVID-19 data to validate the robustness of our methodology. Our strategy to integrate TDA with the clustering algorithm significantly enhanced the performance of sparse portfolios across various performance measures in diverse market scenarios.
Keywords: Portfolio optimization, topological data analysis, clustering techniques, index tracking, Markowitz model, sparse portfolio construction, investment strategies
@@ 1 Introduction
Asset selection and portfolio optimization are critical tasks for financial institutions such as banks, fund management firms, and individual investors as they seek to manage risk and maximize returns in an increasingly competitive business environment and complex market conditions. Over the years, academics and practitioners have actively developed quantitative techniques and models to assess risk and construct efficient portfolios that perform well out-of-sample. In this pursuit, sparse portfolio selection has become a well-known and extensively researched problem in the fields of financial economics and optimization, where sparseness is a desirable property not only from the point of view of transaction costs but also because it allows investors to concentrate their capital on assets that are most likely to generate returns while minimizing exposure to unforeseen risks or market fluctuations. Moreover, it also offers a potential solution to the limitations of traditional mean-variance ([23, 39]) or index-tracking portfolios, as it focuses on a smaller subset of assets to improve portfolio performance. Recently, data-driven approaches ([10, 7, 52]) have been successfully developed for active and adaptive strategies on both index-tracking and mean-variance problems with sparse portfolios. [^1] Despite its popularity, mean-variance analysis has received criticism for its susceptibility to parameter estimates and assumptions. It is observed that small changes in assumed asset returns, volatilities, or correlations lead to substantial effects on the optimization results, ultimately leading to sub-optimal performance when tested against real-world data. In essence, the classic Markowitz portfolio optimization represents an ill-posed inverse problem. Several techniques have been proposed to address this challenge, including regularization techniques imposing sparseness ([8, 48]) and adopting shrinkage estimators of expected returns and covariance matrix ([50, 68, 18])., [^2] In index tracking, the goal is to construct a portfolio of securities that replicates the performance of a market index. The simplest method to achieve this goal is through full replication, which entails holding all the assets in the target market index in the same proportions. While full replication provides perfect tracking performance in a frictionless market, it is expensive, requiring frequent re-balancing of many stocks. It is sometimes not possible due to restrictions on trading illiquid securities or managerial constraints. An alternative approach is partial replication, which aims to mimic the index performance using a small subset of securities.
@@
The problem of sparse portfolio selection can be reformulated as a two-step problem, with the first step of selecting a subset of securities based on their risk and return characteristics and the second step of obtaining the optimal weights. In this direction, clustering methods have emerged as one of the most innovative strategies in portfolio selection in recent years. This technique for classifying data objects into homogeneous groups based on their similarities proved to be effective in constructing sparse portfolios. For instance, [64] studied the comovement of international equity markets using cluster analysis, although their goal was to study the dynamics of this structural comovement and not portfolio optimization. [42] used factor analysis to screen out certain stock markets before creating a Markowitz portfolio. More recently, [51] tested several clustering algorithms, including hierarchical clustering, using correlation coefficients as similarity measure to group stocks and create optimal portfolios. In this paper, we adopt clustering to select a smaller set of securities that provide diversification benefits to the portfolio.
Other methods in the literature for sparse portfolio selection includes regularization techniques (imposing a constraint on the $\ell_{q}$-norm of portfolio weights) (refer [12, 24, 81]) or including a cardinality constraint on the number of securities to be included in the portfolio to promote sparsity and stable allocations ([6, 45, 76, 82]). [12] was the first to integrate an $\ell_{1}$-norm (LASSO) penalty on the vector of portfolio weights in optimization models to regularize the optimization problem and encourage sparse portfolios. [24] demonstrate that norm-constrained global minimum variance portfolio yields improved out-of-sample performance. [32] and [36] utilize $\ell_{1}$ penalty in the context of replicating hedge funds. [41] considered the mean-variance optimization with an advanced regularization technique called weighted Elastic-Net. The $\ell_{1}$ penalty and its variants have also been studied in [48, 22] to obtain sparse portfolios. Recently, [81] considered fraction regularization $\frac{\ell_{1}}{\ell_{2}}$ on the mean–variance model to promote sparse portfolio selection.
@@
We highlight that the clustering based two-step approach has several advantages over the above-mentioned alternative methods. For instance, adding a cardinality constraint significantly increases the combinatorial complexity of the optimization problem, resulting in a mixed-integer quadratic programming problem that is generally NP-hard and computationally intractable. Heuristic methods, such as genetic algorithms, are typically required to find approximate solutions, which may not guarantee global optimality ([72, 79, 30]). Similarly, the regularization technique involves determining the regularization coefficients through cross-validation, which becomes particularly challenging when each security has its regularization parameter; the approach that has been shown to outperform the approach with a common regularization parameter ([41]). Further, these methods tend to result in risk concentration as the resulting portfolio may not be diversified. This can lead to over-exposure to a single industry sector, thus making the portfolio riskier due to vulnerability to a downturn in that sector ([48, 81]). Furthermore, due to local optima and discontinuities in the search space, both the cardinality constraint problem and the $\ell_{q}$-norm constraint problem are difficult to solve from an optimization standpoint. Finally, the regularization technique LASSO suffers from several known shortcomings, such as selecting at random from equally correlated assets ([11]), being biased for large coefficient values ([29, 36]), and being ineffective in the no-short-sale area ([24]). In contrast, clustering can result in a more diversified and sparse portfolio without increasing the computational burden. Moreover, clustering allows for the combination of judgment and quantitative optimization as it is an intuitive approach.
@@
The key to any clustering technique is the distance function that enables it to identify similarities and dissimilarities among the inputs in the data set (time series of stock returns in our case) and allocate them to different clusters. Traditionally, correlation-based clustering has been employed to locate clusters in the portfolio of financial assets, for instance, [57, 51]. However, it is well known that Pearson’s correlation is not a suitable measure of dependence to be used outside the Gaussian (elliptical) distributions, and, in particular, it does not give an accurate indication and understanding of the actual dependence between risk exposures ([28]). Furthermore, even if two assets are perfectly correlated, an investment in one can lead to positive and negative returns in the other. In this paper, we employ a recently introduced approach known as Topological Data Analysis (TDA) to define the similarity between two time series and propose an efficient data-driven approach to sparse portfolio selection.
TDA offers a novel tool to study the shape and qualitative properties of complex datasets which often have non-linear relationships and highly complex dependence structure. The core idea behind TDA is that data has a shape that conveys valuable information. TDA describes the “shape” of the data by extracting topological features, characterized by $k$-dimensional holes, where $k=0,1,2$ represents connected components, holes, and voids, respectively. TDA, by focusing on the topological features of the data, captures information that traditional methods such as correlation may miss. Thus, studying the latent geometric behavior of assets could reveal clusters of assets with similar dynamics. What’s more, a key advantage of TDA is its capability to extract topological features with minimum statistical assumptions, and the only requirement being the computation of distances between various points in the point cloud.
Persistence homology is one of the powerful tools in TDA that provides a comprehensive understanding of the structure of complex data by analyzing its shape across multiple resolutions while simultaneously tracking the changes in its topological features. The information on all topological features with their birth and death is then summarized in a persistence diagram. Notably, persistence diagram, obtained using a point-cloud corresponding to a time-series, can capture various features of the data. For instance, higher fluctuations in time-series often result in more complex topological structures, while structural breaks or different regimes can alter the shape of the data. Further, the higher the concentration (scattering) of the point cloud, and thus the more (less) stable the returns are over the sub-intervals from a topological point of view. By this way, the topological properties of the return data are learnt to capture the variability of returns from the topological point of view.
A comparison among persistence diagrams corresponding to different assets could reveal interesting insights that correlation may miss. While correlation measures linear dependence time-series of two assets, TDA focuses on their qualitative topological properties. TDA based clusters may align with those obtained using correlation, but the former can reveal additional insights particularly in scenarios with non-linear dependencies or structural breaks, a typical behavior observed the financial assets. Thus, clustering persistence diagrams could help identify clusters of assets with similar risk-return properties, leading to more informed portfolios.
Recent developments in TDA have shown significant potential in extracting topological features of data, leading to an improved performance when combined with other models, such as traditional statistical models or machine learning models. The foundation of TDA was laid down by the pioneer contributions of [27] and [85], while an overview paper by [16] turned out to be cornerstone. Thereafter, TDA has been widely applied in various fields, such as time-series clustering and classification ([55, 46]), computer science ([66]), biology ([53]), semi-conductors manufacturing ([47]). TDA only started seeing application in finance very recently. For example, [33] and [35] used TDA to detect crash early warnings, [38] applied persistence landscapes to create portfolios. [80], [31] and [60] integrated TDA and machine learning with persistence diagrams and studied their applications in various fields including finance. [69] studied corporate failure by mapping Altman’s z-score through TDA, while [62] used TDA to detect anomalies in time-series of graphs. Recently, [56] proposed TDA-based distance to measure dependence between stochastic processes and used those distance measures in pair trading. Motivated by the success of TDA, this paper argues that topological features can help identify exciting time series patterns and spatial data clustering. In particular, insights can be gained by examining the essential qualitative characteristics of data sets that are often overlooked by traditional distance-based clustering techniques. To our knowledge, this paper is the first to present a pioneering application of TDA to address the sparse portfolio selection using clustering.
@@
Our paper makes several methodological and empirical contributions that address the aforementioned issues. First, despite the considerable amount of research that has been conducted on applying TDA to study financial time-series, there still needs to be practical implementation and utilization of this method in the field of quantitative finance. [^3] Recently, [37] proposed TDA based sparse index tracking model where the authors adopt the regularization approach to achieve sparseness. We leverage tools from TDA, namely persistence diagrams and landscapes, and design strategies for sparse portfolio selection. By doing so, we establish a link between the growing research on TDA and the most frequently studied portfolio selection problem in the financial economics literature. Thus, we contribute to the emerging literature on applications of TDA to the problem of selecting sparse portfolios.
@@
We conduct an extensive empirical analysis, investigating the performance of the sparse index tracking, MV, and GMV portfolios obtained using the daily returns of the S&P 500 index and its constituents spanning more than a decade from Dec 2009 to July 2019. We further assess the performance of our strategies during the recent COVID-19 Pandemic period from August 2019 to August 2020 to check the robustness. We further utilize the period, September 2020 to August 2022, for comparison purposes, in particular, to other clustering approaches such as $K$-medoids and hierarchical clustering. We also compare the performance of our sparse index tracking model on the benchmark data set for index tracking problems available publicly on the Beasley OR library. Our analysis shows that our TDA-based sparse portfolios outperform the traditional correlation-based sparse portfolios as well as other leading benchmark models in the literature.
@@ 2 Background of TDA
Definition 2.
The $p$-Wasserstein distance [21] between two persistence diagrams ${\mathcal{D}}_{{X}}$ and ${\mathcal{D}}_{{Y}}$ corresponding to data sets $X$ and $Y$ is defined as
@@
Though persistence diagrams contain potentially valuable information about the underlying data set, it is a multi-set which, when equipped with Wasserstein distance, forms an incomplete metric space and hence is not appropriate to apply tools from statistics and machine learning for processing topological features ([14]). A persistence landscape is an alternative tool to convey the information in the persistence diagram. It comprises a sequence of continuous and piece-wise linear functions defined on a re-scaled birth-death coordinate, with the peaks of a persistence landscape representing the significant topological features. Given the birth-death pairs $p_{i}(a_{i},b_{i}),~{}i=1,\ldots,m$, a persistence landscape, illustrated in Figure 3(b), is defined as a sequence of functions $\lambda:\mathbb{N}\times\mathbb{R}\rightarrow[0,\infty)$, denoted as $\lambda(k,t)=\lambda_{k}(t)$. Here, $\lambda_{k}(t)$ represents the $k$-th largest value among the set $\{\Lambda_{p_{i}}(t)\mid i=1,\ldots,m\}$ and we set $\lambda_{k}(x)=0$, ensuring that $\lambda_{k}(t)=0$ for $k>m$ if the $k$-th largest value does not exist. Here, for a fixed feature dimension $f$, we associate a piece-wise linear function $\Lambda_{p}(t):\mathbb{R}\rightarrow[0,\infty)$ with each birth-death pair $p=(a,b)\in{\mathcal{D}}$, where ${\mathcal{D}}$ is the persistence diagram, as follows:
@@
One of the notable advantages of the persistence landscape representation, as compared to persistence diagrams (refer to [14, 13]), is that it forms a subset of the Banach space $L^{p}(\mathbb{N}\times\mathbb{R})$, which comprises sequences $\lambda=(\lambda_{k})_{k\in\mathbb{N}}$. This set possesses an inherent vector space structure and becomes a Banach space when equipped with a norm that measures the degree of “spread out” or “concentration” exhibited by the landscape functions. The norm of a persistence landscape characterizes a point cloud by a single numerical value, and a higher norm value indicates the presence of more significant topological features in the underlying point cloud and is given by
@@
Given that TDA operates on point clouds in multi-dimensional spaces, and our focus in this paper is on time series of stock returns, which lack natural point cloud representations, we must devise a method to obtain a point cloud from a time series. To achieve this, we adopt the standard procedure, called Taken’s embedding ([78]) that embeds a time series $x=\{x_{t},t=1,\ldots,T\}$ to a multi-dimensional representation as follows:
@@ 3 New distance measures
The notion of distance is inherently linked to the characteristics of the underlying data. For instance, the distance between two financial time series may differ from that between two DNA sequences. Nonetheless, distance measure aims to capture some sense of similarity, and often, one can employ the distance measure inherited from the reference space. However, if that is not the case, one must use a more suitable metric. For example, to compare persistence diagrams, Wasserstein distance (or bottleneck distance) is traditionally used due to their stability concerning perturbations of the input [35, 34, 38]. However, considering the time factor while dealing with financial time series is crucial, which the traditional Wasserstein distance fails to capture, as shown in the following toy example.
@@
While the example above is hypothetical, it advocates for a more appropriate distance measure for time-series data. It emphasizes that Wasserstein distance might be beneficial when dealing with observations from a fixed geometrical object where one is trying to identify the structure of that object. Still, it is unsuitable for financial applications where time is essential. Further, the literature has observed that computing the Bottleneck or the Wasserstein distances is computationally demanding as it requires finding point correspondence as observed in [2]. Therefore, we propose new distance measures between persistence diagrams and persistence landscapes that are efficient in terms of time and also consider the time component inherent in time-series data.
@@
where ${\mathcal{D}}_{X-Y}$ represents the persistence diagram corresponding to the point cloud obtained from the series ${x}-{y}$ and ${\mathcal{D}}_{0}$ is the diagram corresponding to a constant series (note that PDs are shift-invariant), containing only the diagonal [59], so that $WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{0})=\left(2^{-p}\sum_{a\in{\mathcal{D}}_{X}}(pers(a))^{p}\right)^{\frac{1}{p}}$, where $pers(a)$ denotes the persistence of $a=(b,d)\in{\mathcal{D}}_{x}$, given by $pers(a)=d-b$, with $b$ and $d$ denoting the birth and death time.
As highlighted above, the traditional Wasserstein distance, $WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})$, has notable limitations. In the above toy-example, we observe that it successfully captures the topological differences between the time-series $x$ and $y$, but misses out to reflect the point-wise similarities as it is sensitive to phase shifts due to its dependence on the positioning of the topological features in $x$ and $y$. Additionally, computing $WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})$ can be computationally intensive as it requires evaluating all bijections from the points in ${\mathcal{D}}_{{X}}$ to the points in ${\mathcal{D}}_{{Y}}$.
The average Wasserstein distance addresses this by utilizing a sliding window approach, dividing the time-series to multiple sub-series that allows to measure the changes in the shape of the point-cloud associated to each sub-series over the time-span of time-series. The robustness inherited in persistent homology against small perturbations makes average Wasserstein distance suitable primarily for analyzing noisy time-series. Further, it also facilitates capturing the abrupt shifts from one regime to other due to minor changes in external factors, called critical transitions, as demonstrated in [3].
On the other hand, the differenced $p$-Wasserstein distance ($DWD_{p}$) aims to capture the point-wise discrepancies in the time-series’ $x$ and $y$ over time. Clearly, a small value of $DWD_{p}$ implies that $x-y$ are close to zero implying strong topological similarity between $x$ and $y$. Moreover, it simplifies the computation of Wasserstein distance as it reduces the complexity by comparing a persistence diagram ${\mathcal{D}}_{X-Y}$ to a constant persistence diagram. Furthermore, as it captures the point-wise similarity between $x$ and $y$, thus, it is more robust to small mis-alignments in the two time-series. This property, along with evaluation of similarity between $x$ and $y$ against a benchmark of zero, make it particularly useful for applications such as detecting anomalies or time-series matching.